QUESTION. AD-AS Consider a theoretical framework where the short-run AS describes, as usual, a positive relationship between the price and output levels. The AD is obtained by combining the following equations:

C = c0 + c1(Y − T)
I = d0 − d2i
T = t0 + t1Y
L = f1Y − f2i
with G = G0 and M = M0.
Starting from a long run equilibrium allocation, assume that autonomous consumption
decreases, ∆c0 = x < 0

a) Represent graphically (AD-AS plot) the short run effect indicating if and how output
and price change

b) Moving from the starting long-run equilibrium to the short-run equilibrium, indicate if
and in which direction the following variables have changed: i interest rate, I invest-
ment, C consumption, T taxation, G public expenditure, w nominal wage
1
Expert's answer
2016-04-12T08:59:05-0400
C = c0 + c1(Y − T), I = d0 − d2i, T = t0 + t1Y, L = f1Y − f2i, with G = G0 and M = M0.
If autonomous consumption decreases, ∆c0 = x < 0, then:
a) Then according to decrease in consumption and in total aggregate demand (AD) the output and price will decrease.
b) Moving from the starting long-run equilibrium to the short-run equilibrium, interest rate (i) will decrease, investment (I) will decrease, consumption (C) will decrease, taxation (T) will decrease, public expenditure (G) will not change and nominal wage (w) will not change too.